Controls Meeting 7-Feb-2013
- Permanent room change to A110
- Josh leaving in 3 weeks, Dave takes over, who will be backup?
- Minutes from previous meeting 24-Jan-2013
- Saga of the EPICS archiver - Hovanes, Elliott, Graham
- Experience with Elog - Elliott, Hovanes, Graham
- Network outage in solenoid control system last week - Josh, Hovanes, Elliott
- how might we determine the PLC is operating properly and controlling all tags?
- Update on gas system - Dave, Josh
- Counting House computer system - Hovanes, Elliott
- file server is central component
- Windows applications - PC's vs terminal server
- start using code management (SVN)
- general (non-whitelisted) web access?
- where to put computers in the hall?
- A-B license servers?
- new printer, Tom's printer, scanner
- UPS power in the hall - Elliott
- EPICS update - Hovanes
- Solenoid test status - Josh, Mark, Scot, Elliott
- MySQL and logging
- SOE, PXI and UPS, Danfysik fan and UPS, QD, EWEB
- Other controls applications
- Use code management for PLC code once file server installed
- FCAL, BCAL, FDC cooling, etc.
- JInventory - Elliott
- barcode readers, label printers and pre-printed labels, conventions, etc.
- Mantis task tracker
- Midas elog
- Experiment Control Supervisor - Elliott
Note permanent room change!
9:30AM Thurs 7-Feb-2013 CC A110
Skipping a meeting due to collaboration meeting
9:30am Thurs 7-Mar-2013
(written after the collaboration meeting based on faulty memory and inadequate notes!)
Present: Elliott W, Dave B, Graham H (forgot to record who else).
Still negotiating with Accelerator, hoping to bring computer cost down from original $55k.
So far so good for solenoid test, but many requested features not implemented. Basic entries w/attachments work, as does autolog entries, comments and updates.
Do NOT muck with network equipment, contact Elliott, Hovanes or David L prior to touching anything. We lost connection to the PLC for a few hours due to a cabling mistake.
Dave will order PLC equipment. He will test PLC and some sensors in his test setup in the ARC.
We are in the process of ordering/borrowing computer equipment and should have a basic system installed by the end of the month. We will purchase remaining equipment in conjunction with the upcomong large CC order (may be able to save a lot). Furniture and consoles are mostly chosen and we hope to order everything by early next month.
UPS Power in the Hall
Final decision by Giles, but it looks like using RFI line filters to filter power from the Counting House UPS and running circuits down to the hall is the way to go. Fernando is happy with this.
Still waiting for Cryo to fix compressors, might go cold in a few weeks. Danfysik power supply to be checked out, including new PLC code (by Dave). SOE almost ready, needs to be checked. QD needs modifications and testing. PXI needs to be tested. Some cables still to be installed. MySQL logging by RSLogix program working well after database mods.